Grade 11General Physics

a motor boat moves at a steady speed of 8 m/s, if the water resistance to the motion of the boat is 2000 N. calculate teh power of the engine.

ranganathan
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
 power is the rate of doing work  =dw/dt
                                          = dFx/dt (work = Fx)
                                            = Fv
power of engine = 2000*8
                         = 16000w
                         = 16kw